OpenStudy (anonymous):

If the volume of a given mass of gas, at atmospheric pressure and 150C, is doubled, the new temperature will be a. 300C b. 450C c. 500C d. 573C

OpenStudy (anonymous):

300C, us the equation PV = nRT for an ideal gas. Pressure remains constant (atmospheric pressure) and the number of moles (n) also remains constant given by a fixed mass of gas. If V doubles so does T

OpenStudy (anonymous):

use*

OpenStudy (anonymous):

And T is in Celsius, not Kelvin?

OpenStudy (anonymous):

oh sorry, T has to be in Kelvin, so we get (150+273)*2 in K which is 846 K and hence 573C

OpenStudy (anonymous):

Are you sure it's in Kelvin?

OpenStudy (anonymous):

Yes.In the ideal gas equation,temperature values have to be plugged in to it in Kelvin

OpenStudy (anonymous):

If I were to ask what temperature is twice as hot as 150 C, the answer would not be 300 C. Temperature is the measure of heat, but the Celsius scale is displaced by 273 degrees from the true thermodynamic temperature as measured in Kelvin. The Kelvin scale is the absolute scale, beginning at absolute zero, (the lowest possible temperature). Absolute zero is 0 K on the Kelvin scale, and -273 C on the Celsius Scale. If I asked what temperature is twice as hot as 150 K, the answer would be 300 K. So to answer the original question, we would have to convert 150 C to kelvin , then double it, then convert back to Celsius. Thus we have 150 C = (273+150) K = 423 K. Now double it and we get 846 K. Now convert back to Celsius 826 K = (846-273) = 573 C. All temperatures when using any thermodynamic equation such as the ideal gas equation must be converted to Kelvin, otherwise wrong answers are obtained.

OpenStudy (anonymous):

The answer is 573C... Use Charle's Law. V1/T1=V2/T2. Replace V1 with V, T1 with (273+150=) 423 and V2 by 2*V.... and solve the equation... U wil get 846... Subtract 273 n u'l get 573C...